you have a rectangle that 10 m wide and 24 m long
how long is the diagonal?

Answer by jim_thompson5910(35256)

a^2 + b^2 = c^2

10^2 + 24^2 = c^2

100 + 576 = c^2

676 = c^2

c^2 = 676

c = sqrt(676)

c = 26


So the diagonal is 26 m long